Scores on a recent national statistics exam were normally distributed with a mean of 72 and a standard deviation of 10 . What is the probability that a randomly selected exam will have a score between 75 and 80 ?
Answer #1

Solution :

Given that ,

mean = = 72

standard deviation = = 10

P(75< x < 80) = P[(75-72) /10 < (x - ) / < (80-72) /10 )]

= P(0.3 < Z < 0.8)

= P(Z < 0.8) - P(Z <0.3 )

Using z table   

= 0.7881-0.6179

=0.1702